Five times Jeff's age plus three times Karen's are equals 92. Karen's age is 4 more than Jeff's age. How old is Karen?
Answer by VFBundy(438)

Karen's age = k
Jeff's age = j

5j + 3k = 92
k = j + 4

5j + 3k = 92
-j + k = 4

Multiply second equation by 5:

5j + 3k = 92
-5j + 5k = 20

Add both equations together and solve for k:

8k = 112

k = 14

Karen's age = k = 14